OpenStudy (anonymous):

compute d^2/dx^2 at the point (4,2) x^2+y^2=20

OpenStudy (amistre64):

is that our f(x) sitting there?

OpenStudy (amistre64):

2x x' + 2y y' = 0 2x + 2y y' = 0 y' = -x/y

OpenStudy (amistre64):

y'' = -1/y + y'x/y^2

OpenStudy (amistre64):

\[y'' = -\frac{1}{y} + \frac{{-x^2}}{y^3}\]

OpenStudy (amistre64):

(x=4, y=2) y'' = -1/2 - 16/8

OpenStudy (amistre64):

y'' = -32/8 = -4 perhaps?

OpenStudy (amistre64):

thats wrong

OpenStudy (anonymous):

why not 5/2?

OpenStudy (amistre64):

-4/8 - 16/8 = -20/8 y'' = -5/2

OpenStudy (amistre64):

i was gettin there, my head on holds so much at once lol

OpenStudy (anonymous):

cool, thanks!

OpenStudy (amistre64):

i was wondering if i could have done that from this step like this: 2x + 2y y' = 0 2 + 2y' + 2y y'' = 0 y'' = -2 -2(-x/y) -1 -(-x/y) -1 -(-2) ---------- = -------- = ------- = doesnt look like it 2y y 2

OpenStudy (anonymous):

that looks more complicated, do you happen to also know about definite integrals?

OpenStudy (amistre64):

i can do them with somewhat of skill and mostly luck ;)

OpenStudy (amistre64):

definite integrals are integrating within an interval

OpenStudy (anonymous):

lol okay, let's see how lucky you are, I'm about to post one =D

OpenStudy (amistre64):

yay!!

OpenStudy (anonymous):

okay, for some reason it won't let me ask a question like normal

OpenStudy (amistre64):

refresh your browser

OpenStudy (amistre64):

f5